OpenStudy (danise):

Solve the equation using he Quadratic Formula x^2 - 7x -3=0

OpenStudy (kidrah69):

list your a b c values :)

OpenStudy (danise):

Ok what else I did that ?

OpenStudy (kidrah69):

well then we plug them in

OpenStudy (kidrah69):

First tell me what they are

OpenStudy (kidrah69):

\[ax^2+bx+c\]

OpenStudy (kidrah69):

Once you have the values you plug them in this formula \[x=\frac{-b ± \sqrt(b^2 - 4ac) }{ 2a }\]

OpenStudy (danise):

A: x B:-7 C:-3

OpenStudy (kidrah69):

\[\rm x=\frac{-b ± \sqrt{(b^2 - 4ac)} }{ 2a }\] now plug it in

OpenStudy (kidrah69):

\[\rm x=\frac{-(-7) ± \sqrt{(-7)^2 - 4(1)(-3)} }{ 2(1) }\]

OpenStudy (kidrah69):

From there all you do is solve

OpenStudy (danise):

Okay do I multiply or divide

OpenStudy (kidrah69):

fist solve whats in the square root but do not square root it just do it as an equation

OpenStudy (danise):

ok let try to solve it

OpenStudy (danise):

is the answer -1

OpenStudy (danise):

then I divide by the bottom

OpenStudy (kidrah69):

no its not

OpenStudy (kidrah69):

\[\rm x=\frac{7 ± \sqrt{61} }{ 2 }\] now to solve it all you do is set up 2 equations \[\rm x=\frac{7 +\sqrt{61} }{ 2 }\] \[\rm x=\frac{7 - \sqrt{61} }{ 2 }\]

OpenStudy (kidrah69):

Then you solve.

OpenStudy (danise):

do i square root 61?

OpenStudy (kidrah69):

yes

OpenStudy (kidrah69):

http://www.math.com/students/calculators/source/quadratic.htm i suggest using that instead of trying to calculate it by hand

OpenStudy (danise):

ok thank you

OpenStudy (kidrah69):

No problem.